# Wasserstein Flow Matching: Generative modeling over families of distributions

### v1

@@ A.2 Entropic transport map: An out-of-sample estimator
This estimator was initially to provide statistical approximations to the optimal transport map $T_{\star}^{\mu\to\nu}$ on the basis of samples; see Pooladian and Niles-Weed, (2021); Pooladian et al., 2023b ; Pooladian et al., (2022). Note that $\hat{T}_{\varepsilon}(x)$ can be interpreted as the conditional expectation of the plan $\bm{P}^{\star}$ conditioned on out-of-sample inputs $x\in\mathbb{R}^{d}$, which is well-defined due to the relations above. Finally, we stress that this estimator can be adapted to settings where the point-clouds $\mu$ and $\nu$ not only have different numbers of points, but also non-uniform weights. As this estimator is also a by-product of Sinkhorn’s algorithm, it is also scalable and GPU-friendly.
Appendix B Multisample Wasserstein Flow Matching
@@
Since optimal transport can be applied on the Wasserstein manifold itself, both WFM and BW-FM can be seamlessly integrated with the multisample FM (MS-FM) framework (Pooladian et al., 2023a, ; Tong et al.,, 2023). The core technique behind MS-FM is to use OT to match minibatches from source and target measures during training, rather than relying on random pairings. This has shown to improve learned flows while requiring fewer function evaluations to synthesize new samples. Applying MS-FM requires computing the pairwise distance matrix between source and target batch samples, denoted from $i\in\{1,\ldots,\texttt{Bsz}\}$. In the BW-FM setting, given two sets of Gaussians $\{(a_{i},A_{i})\}_{i=1}^{\texttt{Bsz}}$ and $\{(b_{i},B_{i})\}_{i=1}^{\texttt{Bsz}}$, their Frechét ($W^{2}_{2}$) distance matrix is:
@@
We then use entropic OT to approximately solve the assignment problem on $C$ and compute a transport matrix. This is the converted into a one-to-one assignment matrix via rounded matching (Section A.1), ensuring the entire batch is used in training.
For WFM point-clouds, applying MS requires computing pairwise OT distance between all source and target samples within a minibatch. For large point-clouds, this is exorbitantly expensive, even with Sinkhorn iterations. For an efficient approximate, here too we rely on the Frechét distance, computed between empirical means and covariances of each point-cloud. Computation of the Frechét distance is markedly less resource-intensive than entropic OT, yet is notably correlated with EMD values (see Table 1 in (Haviv et al., 2024b, )).
Appendix C Sampling from source measure
@@
To achieve high-quality generation of point-clouds, it is essential that the initial (source) distribution be diverse, rather than collapsed and degenerate. Indeed, while it is alluring to produce noisy point-clouds by sampling points from a single base distribution, i.e. $X=\{x_{i}\}_{i=1}^{n},\,x_{i}\sim\mathcal{N}(0,I_{d})$, as $n$ grows, the Wasserstein distance between instances goes to 00. To alleviate this, we draw point-clouds from multivariate Gaussians with a stochastic covariance:
@@
Figure S1: Conditional BW-FM applied to single-cell RNA sequencing data of immune response to COVID-19. Large scale single-cell atlases are commonly grouped into highly dedicated clusters called MetaCells (Persad et al.,, 2023). In this application, BW-FM is conditioned on cell state and trained to generate means and covariances of gene expression, focusing on the top 32 principal components, derived from aggregated cells. The model achieves high-quality sample generation, as evidenced by a label accuracy of 93.13%.
Appendix D Neural architecture & training
D.1 BW-FM on Gaussians
@@
The goal of BW-FM is to train a neural network to match the (Riemannian) time-derivative along the BW geodesics between Gaussians. The model employs a standard, fully connected neural network which takes as input concatenated values of $(m_{t},\Sigma_{t},t)$ based on the McCann interpolation formula from Section 2.3.1. Since the covariance matrix is symmetric, only its lower-diagonal values are used, flattened into a vector of length $d(d+1)/2$. Time values are converted to Fourier features, an approach inspired by positional encodings in transformer literature (Vaswani,, 2017). To streamline training, two separate networks are employed: one to match the time derivative of the mean $\dot{m}_{t}$ and another for the time derivative of the covariance matrix $\dot{\Sigma}^{{\rm{BW}}}_{t}$. The BW tangent norm is used as the loss function for training these networks.
@@
By default, all models use a $6$-layer neural network using relu non-linearity, with $1024$ neurons per layer, applying skip connections and layer-norm Ba, (2016). Training is performed for $100,000$ gradient descent steps using the Adam optimizer (Kingma,, 2014) with an exponential learning rate decay of $0.97$ every $1000$ steps and batch size of $128$.
Figure S2: Spatial arrangement of microglia in the motor cortex. Bures-Wasserstein distance based 2D UMAP visualization of real microglia and BW-FM synthesized niches, colored by their first diffusion component (DC). This DC corresponds to the cortical depth of the microglia across the MERFISH slices.
D.2 WFM on point-clouds
@@
WFM is designed to estimate the optimal transport (OT) map for a given pair of interpolate point-cloud and time $(\bm{X}_{t},t)$. Here too the time component $t$ is first converted into Fourier features. The model’s architechture begins with an embedding layer, followed by a series of alternating multihead attention and fully-connected layers. Skip connections and layer-norma are applied after each operation. The final layer projects the embeddings back to $X$’s original space using a dense layer with zero initialization. The model is trained by minimizing the squared distance between the predicted and true OT maps.
@@
By default, the entropic OT map is constructed with regularization weight of $\varepsilon=0.002$ and $200$ Sinkhorn iterations, which we found to be sufficient for convergence. Whenever the dataset consists of uniformly sized point-clouds, we use rounded matching (Section A.1), otherwise we apply the out-of-sample estimator (Section A.2) which can calculate maps between point-clouds with different sizes. The transformer network is composed to $6$ multi-head attention blocks, with an embedding dimension of $512$ and $4$ heads. The model is trained for $500,000$ steps, using the Adam optimizer with exponential learning rate decay and batch size of $64$.
Appendix E Experiment Details
E.1 Spatial Transcriptomics

### v2

@@ A.2 Entropic transport map: An out-of-sample estimator
This estimator was initially to provide statistical approximations to the optimal transport map $T_{\star}^{\mu\to\nu}$ on the basis of samples; see Pooladian & Niles-Weed (2021); Rigollet & Stromme (2022); Pooladian et al. (2023b, 2022). Note that $\hat{T}_{\varepsilon}(x)$ can be interpreted as the conditional expectation of the plan $\bm{P}^{\star}$ conditioned on out-of-sample inputs $x\in\mathbb{R}^{d}$, which is well-defined due to the relations above. Finally, we stress that this estimator can be adapted to settings where the point-clouds $\mu$ and $\nu$ not only have different numbers of points, but also non-uniform weights. As this estimator is also a by-product of Sinkhorn’s algorithm, it is also scalable and GPU-friendly.
A.3 On the (statistical) approximations of geodesics
We briefly collect a basic results pertaining to the (statistical) approximation of optimal transport paths. This bound shows that the error grows along the trajectory, but is limited by the overall distance of the maps.
Proposition A.1.
Let $\mu,\nu$ be two probability measures and suppose $\mu$ has a density, and let $T^{\star}$ be the optimal transport map from $\mu$ to $\nu$. Let $\hat{T}$ be an estimator to the optimal transport map, defined with respect to data $X_{1},\ldots,X_{n}\sim\mu$ and $Y_{1},\ldots,Y_{n}\sim\nu$. Then for $t\in[0,1]$
$\mathbb{E}[W_{2}^{2}(\rho_{t},\hat{\rho}_{t})]\leq t^{2}\mathbb{E}\|\hat{T}-T^{\star}\|^{2}_{L^{2}(\mu)}\,,$
where the outer expectation is taken with respect to the data, and we define
$\rho_{t}\coloneqq((1-t){\rm{id}}+tT^{\star})_{\sharp}\mu\,,\quad\hat{\rho}_{t}\coloneqq((1-t){\rm{id}}+t\hat{T})_{\sharp}\mu\,$
Proof.
The result follows immediately from a standard coupling argument to obtain the linearized Wasserstein distance (Wang et al., 2010; Panaretos & Zemel, 2020)
$W_{2}^{2}(\rho_{t},\hat{\rho}_{t})$ $\leq\|((1-t){\rm{id}}+tT^{\star})-((1-t){\rm{id}}+t\hat{T})\|^{2}_{L^{2}(\mu)}=t^{2}\|\hat{T}-T^{\star}\|^{2}_{L^{2}(\mu)}\,.\qed$
The entropic Brenier map is one particular estimator. We note two key properties of this map; see Pooladian & Niles-Weed (2021) for in-depth discussions.
Theorem A.2.
Suppose $\mu,\nu$ have density bounded above and below, and that the optimal transport map between them, denoted $T^{\star}$, is such that $(T^{\star})^{-1}$ is at least twice differentiable and there exists $\lambda,\Lambda>0$ such that
$\lambda I\preceq DT^{\star}\preceq\Lambda I\,.$
Then, when estimated from $n$ samples from $\mu$ and $n$ samples from $\mu$, the entropic Brenier map has the following error
$\mathbb{E}\|\hat{T}_{\varepsilon}-T^{\star}\|^{2}_{L^{2}(\mu)}\lesssim n^{-1/2}\log(n)\varepsilon^{-d/2-1}+\varepsilon^{2}\,,$ (14)
where we suppress constants that depend on our assumptions. Performing a bias-variance trade-off in the regularization parameter, one obtains $\varepsilon=\varepsilon(n)\asymp n^{-1/(d+4)}$ and the total error becomes
$\mathbb{E}\|\hat{T}_{\varepsilon}-T^{\star}\|^{2}_{L^{2}(\mu)}\lesssim_{\log(n)}n^{-2/(d+4)}\,.$
We emphasize that the assumptions in Theorem A.2 are standard in the literature (Hütter & Rigollet, 2021; Deb et al., 2021; Muzellec et al., 2021; Divol et al., 2022). While the rate scales exponentially poorly with the dimension, we stress that existing lower bounds of estimation (see Hütter & Rigollet (2021)) also suffer from the curse of dimensionality, which is unavoidable for this task. Combining these two results, we can compare the geodesic given by the OT map, and the one induced by using the entropic map, where we write
$\hat{\rho}_{t}^{\varepsilon}\coloneqq((1-t)\rm{id}+t\hat{T}_{\varepsilon})_{\sharp}\mu\,.$
Corollary A.3.
Consider the same setting as Theorem A.2. Then the geodesic given by the estimated entropic Brenier map, denoted by $\hat{\rho}_{t}^{\varepsilon}$, on the basis of $n$ samples and $\varepsilon\asymp n^{-1/(d+4)}$, is close to the true geodesic with the following error
$\mathbb{E}[W_{2}^{2}(\rho_{t},\hat{\rho}_{t}^{\varepsilon})]\lesssim t^{2}n^{-2/(d+4)}\,.$
Appendix B Derivation of the WFM objective
In this section, we give validity to the WFM for optimization purposes, and our choice of curves. For instance, recall the original Flow Matching objective (Lipman et al., 2022)
$\mathcal{L}_{\rm{FM}}(\theta)\coloneqq\int_{0}^{1}\mathbb{E}_{X_{t}\sim p_{t}}\|f_{\theta}(X_{t},t)-u_{t}(X_{t})\|^{2}\,\mathrm{d}t\,,$ (15)
where $f_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ is a neural network, and $(p_{t},u_{t})_{t\in[0,1]}$ are a density-vector field pair that satisfy the continuity equation between two distribution $\mu$ and $\nu$.
$\partial p_{t}+\nabla\cdot(p_{t}u_{t})=0\,,\quad p_{0}=\mu\,,p_{1}=\nu\,.$
Note that, implicit in $\mathcal{L}_{\rm{FM}}(\theta)$ is the endpoint constraints $\mu$ and $\nu$. Now, we average over possibly choices of $\mu\sim\mathfrak{p}_{0}$ and $\nu\sim\mathfrak{p}_{1}$, resulting in
$\int_{0}^{1}\iint\mathbb{E}_{X_{t}\sim p_{t}}\|f_{\theta}(X_{t},t)-u_{t}(X_{t})\|^{2}\,\mathrm{d}\mathfrak{p}_{0}(\mu)\,\mathrm{d}\mathfrak{p}_{1}(\nu)\,\mathrm{d}t\,.$ (16)
As a particular case, take $(p_{t},u_{t})\leftarrow(\mu_{t},v_{t})$, where the first argument is the McCann interpolation between $\mu$ and $\nu$, and $v_{t}$ is the optimal velocity field, which is a function of the optimal transport map from $\mu$ to $\nu$ (recall Section 2.3). This yields our final objective (11), which we recall here for convenience
$\mathcal{L}_{\rm{WFM}}(\theta)\coloneqq\int_{0}^{1}\iint\mathbb{E}_{X_{t}\sim\mu_{t}}\|f_{\theta}(X_{t},t)-v_{t}(X_{t})\|^{2}\,\mathrm{d}\mathfrak{p}_{0}(\mu)\,\mathrm{d}\mathfrak{p}_{1}(\nu)\,\mathrm{d}t\,.$ (17)
When $\mathfrak{p}_{0},\mathfrak{p}_{1}$ are distributions over Gaussians, we have closed-form expressions for all objects of interest. When $\mathfrak{p}_{0},\mathfrak{p}_{1}$ are distributions over general distributions, we approximate the geodesics based on point-cloud samples using entropic Brenier maps and their respective interpolations. We emphasize that the rounded-matching we employ (see Section A.1) is also a valid curve.
B.1 Validity of Conditional Flows
The key idea behind Flow Matching is that the (tractable) conditional velocity probability paths come together to form correct marginal velocities and paths from source distribution to target. In Riemannian Flow Matching, given a source and target $\mathfrak{p}_{1},\mathfrak{p}_{0}\in\mathcal{P}(\mathcal{M})$ distributions over Riemannian manifold $\mathcal{M}$, the conditional velocity field from $\mathfrak{p}_{0}$ to a sample $\nu\sim\mathfrak{p}_{1}$ is generated by flowing from each point $\mu\sim\mathfrak{p}_{0}$ towards $\nu$ using the geodesic path between them. This produces the conditional probability path $\mathfrak{p}_{t}(\cdot|\nu)$ generated by the vector field $\mathfrak{u}_{t}(\cdot|\nu)$, which obey the continuity equation for Riemannian manifolds:
${\partial_{t}}\mathfrak{p}_{t}(\cdot|\nu)+\text{div}_{\mathcal{M}}(\mathfrak{p}_{t}(\cdot|\nu)\mathfrak{u}_{t}(\cdot|\nu))=0$ (18)
where $\text{div}_{\mathcal{M}}$ is the Riemannian divergence operator. Following Theorem 1 in Lipman et al. (2022) and equation 7 in Chen & Lipman (2023), the marginal vector field:
$\mathfrak{u}_{t}(\mu_{t})=\int_{\mathcal{M}}\mathfrak{u}_{t}(\mu_{t}|\nu)\frac{\mathfrak{p}_{t}(\mu_{t}|\nu)}{\mathfrak{p}_{t}(\mu_{t})}\,\mathrm{d}\mathfrak{p}_{1}(\nu)$ (19)
generates the marginal probability path:
$\mathfrak{p}_{t}(\mu_{t})=\int_{\mathcal{M}}\mathfrak{p}_{t}(\mu_{t}|\nu)\,\mathrm{d}\mathfrak{p}_{1}(\nu)$ (20)
as $\mathfrak{p}_{t}(\mu_{t})$ and $\mathfrak{u}_{t}(\mu_{t})$ together obey the continuity equation. The proof requires regularity constraints on $\mathfrak{p}_{t}(\mu_{t}|\nu)$ and $\mathfrak{u}_{t}(\mu_{t}|\nu)$, which we assume as in Chen & Lipman (2023). A withstanding requirement that we need is that a unique optimal transport map exists between any two measures $(\mu,\nu)\in\mathfrak{p}_{0}\times\mathfrak{p}_{1}$. This can be ensured if, for example, all measures in $\mathfrak{p}_{0}$ have a density, which is assumed throughout the text. Positivity of the path $\mu_{t}$ can be ensured through a myriad of conditions (e.g., source and target measures have positive density over the same support). In the explicit Gaussian-to-Gaussian case (i.e., the Bures–Wasserstein manifold), this assumption is trivially satisfied as the manifold is a genuine finite-dimensional Riemannian manifold and the Riemannian machinery of Chen & Lipman (2023) goes through with no modifications. We carry out the rest of our computations for general distributions over the Wasserstein space, and hence the computations are non-rigorous. For a fully rigorous treatment, see Emami & Pass (2025).
Conditional flows and paths for WFM
To continue, we require the following proposition. Recall that a pre-metric is a function ${\rm{d}}:\mathcal{M}\times\mathcal{M}\to\mathbb{R}$ that satisfies the following three requirements
  1. 1.
    ${\rm{d}}(x,y)\geq 0$ for all $x,y\in\mathcal{M}$,
  2. 2.
    ${\rm{d}}(x,y)=0$ if and only if $x=y$,
  3. 3.
    $\nabla_{x}{\rm{d}}(x,y)\neq 0$ if and only if $x\neq y$.
Lemma B.1.
The $2$-Wasserstein distance defines a valid pre-metric over $\mathcal{P}_{2,\rm{ac}}(\mathbb{R}^{d})$.
Proof.
As the $2$-Wasserstein distance is a metric (Villani, 2009; Santambrogio, 2015), the first and second conditions of a pre-metric are trivially satisfied. For the third condition, we note that for $\mu,\nu\in\mathcal{P}_{2,\rm{ac}}(\mathbb{R}^{d})$
$\nabla\tfrac{1}{2}W_{2}^{2}(\cdot,\nu)(\mu)=-\log_{\mu}(\nu)=-2(T^{\mu\to\nu}-\rm{id})$ (21)
For a proof of this computation, see Altschuler et al. (2021) or Zemel & Panaretos (2019). By chain-rule, we have that
$\nabla\tfrac{1}{2}W_{2}^{2}(\cdot,\nu)(\mu)=W_{2}(\mu,\nu)\nabla W_{2}(\cdot,\nu)(\mu)=-\log_{\mu}(\nu)$
This latter quantity is only zero if and only if $\mu=\nu$, and thus the proof is complete. ∎
We now recall how to construct conditional vector fields as a function of pre-metrics ${\rm{d}}$, as described by Eq. (13) by (Chen & Lipman, 2023). The construction is the following
${u}_{t}(x_{t}|y)$ $=(\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}\ln(1-t)){\rm{d}}(x_{t},y)\nabla_{x}{\rm{d}}(x_{t},y)/\|\nabla_{x}{\rm{d}}(x_{t},y)\|^{2}_{g(x_{t})}\,,$ (22)
where $g$ is the metric on the tangent space at $x_{t}$, and $x_{t}$ is a geodesic in the form
$x_{t}=\exp_{x_{0}}((1-t)\log_{x_{0}}(x_{1}))\,.$
Note that our geodesics are precisely of the form
$\mu_{t}=\exp_{\mu}((1-t)\log_{\mu}(\nu))\,,$
where we recall $v\mapsto\exp_{\mu}(v)=({\rm{id}}+v)_{\sharp}\mu$ and $\nu\mapsto\log_{\mu}(\nu)=T^{\mu\to\nu}-{\rm{id}}$. We now specialize these calculations to the case where the pre-metric is the $2$-Wasserstein distance.
Lemma B.2.
Fix $\nu\sim\mathfrak{p}_{1}$. Let $\mu_{t}\mapsto\mathfrak{u}_{t}(\mu_{t}|\nu)$ denote the conditional vector field defined using Equation 22 with ${\rm{d}}$ replaced by the $2$-Wasserstein distance, evaluated at a geodesic $\mu_{t}$ between any $\mu\sim\mathfrak{p}_{0}$ and the fixed $\nu$. Then it holds that $\mathfrak{u}_{t}(\mu_{t}|\nu)=v_{t}$, where $v_{t}$ is given by Equation 7 between $\mu$ and $\nu$.
Proof.
We compute, using Equation 21,
$\mathfrak{u}_{t}(\mu_{t}|\nu)$ $=(\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}\ln(1-t))W_{2}(\mu_{t},\nu)\nabla W_{2}(\mu_{t},\nu)/\|\nabla W_{2}(\mu_{t},\nu)\|^{2}_{L^{2}(\mu_{t})}$
$=\frac{1}{1-t}(T^{\mu_{t}\to\nu}-{\rm{id}})$
$=\frac{-1}{1-t}(T^{\mu\to\nu}\circ(T_{t})^{-1}-{\rm{id}})$
$=\frac{1}{1-t}(T^{\mu\to\nu}-T_{t})\circ(T_{t})^{-1}\,,$
where we make use of the fact that $T^{\mu_{t}\to\nu}=T^{\mu\to\nu}\circ(T^{\mu_{t}\to\nu})^{-1}$ (see e.g., Section 2 of Carlen & Gangbo (2003)), and that $T_{t}=(1-t){\rm{id}}+t{T^{\mu\to\nu}}$. Simplifying, we obtain
$\mathfrak{u}_{t}(\mu_{t}|\nu)=\frac{1}{1-t}(T^{\mu\to\nu}-(1-t){\rm{id}}-tT^{\mu\to\nu})\circ(T_{t})^{-1}=(T^{\mu\to\nu}-{\rm{id}})\circ(T_{t})^{-1}\,,$ (23)
which is precisely the optimal transport vector field $v_{t}$ stated in (7). ∎
Together, we arrive at the following corollary, which validates the probability paths generated by our approach.
Corollary B.3.
The conditional flow $\mathfrak{p}_{t}(\cdot|\nu)$ generated by the conditional vector fields defined by (23) satisfy $\mathfrak{p}_{1}(\cdot|\nu)=\nu$.
Proof.
This follows immediately from Theorem 3.1 of Chen & Lipman (2023) by passing through the pre-metric construction above. ∎
Appendix C Multisample Wasserstein Flow Matching
@@
Since optimal transport can be applied on the Wasserstein manifold itself, both WFM and BW-FM can be seamlessly integrated with the multisample FM (MS-FM) framework (Pooladian et al., 2023a; Tong et al., 2023). The core technique behind MS-FM is to use OT to match minibatches from source and target measures during training, rather than relying on random pairings. This has shown to improve learned flows while requiring fewer function evaluations to synthesize new samples. Applying MS-FM requires computing the pairwise distance matrix between source and target batch samples, denoted from $i\in\{1,\ldots,\texttt{Bsz}\}$. In the BW-FM setting, given two sets of Gaussians $\{(a_{i},A_{i})\}_{i=1}^{\texttt{Bsz}}$ and $\{(b_{i},B_{i})\}_{i=1}^{\texttt{Bsz}}$, their Frechét ($W^{2}_{2}$) distance matrix is:
@@
We then use entropic OT to approximately solve the assignment problem on $C$ and compute a transport matrix. This is the converted into a one-to-one assignment matrix via rounded matching (see Section A.1), ensuring the entire batch is used in training.
For general WFM, applying MS requires computing pairwise OT distance between all source and target samples within a minibatch. For large point-cloud samples, this is exorbitantly expensive, even with Sinkhorn iterations. For an efficient approximate, here too we rely on the Frechét distance, computed between empirical means and covariances of each point-cloud. Computation of the Frechét distance is markedly less resource-intensive than entropic OT, yet is notably correlated with EMD values (see Table 1 in Haviv et al. (2024b)).
Appendix D Choice of source measure
@@
To achieve high-quality generation of general distributions, it is essential that the initial (source) measure be diverse, rather than collapsed and degenerate. Indeed, while it is alluring to produce noise point-clouds by sampling particles from a single base distribution, i.e. $X=\{x_{i}\}_{i=1}^{n},\,x_{i}\sim\mathcal{N}(0,I_{d})$, as $n$ grows, the Wasserstein distance between empirical distributions goes to 00. To alleviate this, we draw point-clouds from multivariate Gaussians with a stochastic covariance:
@@
Figure S1: When the number of training examples is too few (in comparison to Figure 2), all methods collapse on the training data, though only the Riemannian instantiation of BW-FM captures the covariances perfectly.
Appendix E Neural architecture & training
E.1 BW-FM on Gaussians
@@
The goal of BW-FM is to train a neural network to match the (Riemannian) time-derivative along the BW geodesics between Gaussians. The model employs a standard, fully connected neural network which takes as input concatenated values of $(m_{t},\Sigma_{t},t)$ based on the McCann interpolation formula from Section 2.3.1. Since the covariance matrix is symmetric, only its lower-diagonal values are used, flattened into a vector of length $d(d+1)/2$. Time values are converted to Fourier features, an approach inspired by positional encodings in transformer literature (Vaswani, 2017). To streamline training, two separate networks are employed: one to match the time derivative of the mean $\dot{m}_{t}$ and another for the time derivative of the covariance matrix $\dot{\Sigma}^{{\rm{BW}}}_{t}$. The BW tangent norm is used as the loss function for training these networks.
@@
By default, all models use a $6$-layer neural network using relu non-linearity, with $1024$ neurons per layer, applying skip connections and layer-norm (Ba, 2016). Training is performed for $100,000$ gradient descent steps using the Adam optimizer (Kingma, 2014) with an exponential learning rate decay of $0.97$ every $1000$ steps and batch size of $128$.
E.2 WFM on general distributions
@@
WFM estimates the optimal transport (OT) map for a given pair of interpolate point-cloud and time $(\bm{X}_{t},t)$. Here too the time component $t$ is first converted into Fourier features. The model’s architecture begins with an embedding layer, followed by a series of alternating multi-head attention and fully-connected layers. Skip connections and layer-norm are applied after each operation. The final layer projects the embeddings back to $X$’s original space using a dense layer with zero initialization. The model is trained by minimizing the squared distance between the predicted and true OT maps.
@@
By default, the entropic OT map is constructed with regularization weight of $\varepsilon=0.002$ and $200$ Sinkhorn iterations, which we found to be sufficient for convergence. Whenever the dataset consists of uniformly sized point-cloud realizations, we use rounded matching (see Section A.1), otherwise we apply the out-of-sample estimator (see Section A.2) which can calculate maps between point-clouds with different sizes. The transformer network is composed to $6$ multi-head attention blocks, with an embedding dimension of $512$ and $4$ heads. Our model is optimizer with Adam (Kingma, 2014) using an exponential learning rate decay and batch size of $64$.
Figure S2: Comparing real and generated cellular microenvironments in embryonic gut tube development. Left: UMAP embedding of BW-FM generated (left panel) and observed (middle panel) microenvironments based on Bures-Wasserstein distance, colored by developing organ type. Right: Spatial arrangement of cells in the seqFISH embryo showing physical organization of organ development.
WFM relies on JAX and OTT-JAX (Bradbury et al., 2021; Cuturi et al., 2022) and enjoys seamless optimization via end-to-end just-in-time compilation. For the ShapeNet experiments (see Table 3), the model is trained for 500,000 training steps, totaling to about 3 days of trainings of a single A100 GPU. All other experiments (see Table 4) trained for 100,000 steps, requiring around 3-4 hours of GPU use. We note the Transformer’s forward and backwards pass was the most significant source of computational overhead, as opposed to the Sinkhorn based approximation of OT maps. The computational complexity of both components is quadratic with point-cloud size, which limits the scalability of WFM. Making Transformers simpler to optimize and accelerating OT computations are both active areas within ML research (Amos et al., 2022; Scetbon et al., 2021; Zhang et al., 2022), and future iterations of WFM can incorporate solution from those spaces.
Appendix F Experiment Details
F.1 Spatial Transcriptomics